OpenStudy (anonymous):

who ever gives me the best answer gets a medal What is the solution to r over three is equal to one over three? one over twelve twelve over one three over four four over three

OpenStudy (anonymous):

r=1

jimthompson5910 (jim_thompson5910):

"r over three is equal to one over three" is the equation \[\Large \frac{r}{3} = \frac{1}{3}\] Cross multiply and solve for r to get: \[\Large \frac{r}{3} = \frac{1}{3}\] \[\Large 3*r = 1*3\] \[\Large 3r = 3\] \[\Large r = \frac{3}{3}\] \[\Large r = 1\]

OpenStudy (whpalmer4):

Or, just look at it:\[\frac{r}{3} = \frac{1}{3}\]\[r=1\]if the two are going to be equal, right?

jimthompson5910 (jim_thompson5910):

Because none of your answer choices are '1', I'm assuming there's a typo somewhere.

OpenStudy (anonymous):

guys pick one over twelve twelve over one three over four four over three one of theses are the answers

jimthompson5910 (jim_thompson5910):

draw out the equation for us please

OpenStudy (anonymous):

1/12 or 12/1 or 3/4 or 4/3

jimthompson5910 (jim_thompson5910):

the original equation is r/3 = 1/3 right? or no?

jimthompson5910 (jim_thompson5910):

there has to be a typo somewhere, so I'd ask your teacher about it